# AIR-Bench: Automated Heterogeneous Information Retrieval Benchmark

### v1

@@ 4 Diversity Analysis
multilingual-e5-small Wang et al. (2024) 118M 42.253 11 47.989 15 28.579 16
simlm-base-msmarco-finetuned Wang et al. (2022a) 110M 41.675 12 48.102 14 30.548 9
jina-embeddings-v2-base-en Günther et al. (2023) 137M 39.887 13 51.112 12 29.560 14
jina-embeddings-v3 Sturua et al. (2024) 572M 39.787 14 51.098 13 30.297 12
bge-m3 Chen et al. (2024b) 568M 39.565 15 54.404 6 33.286 3
contriever-msmarco Izacard et al. (2022) 109M 36.570 16 47.127 16 29.231 15
msmarco-roberta-base-ance-firstp Xiong et al. (2021) 125M 33.637 17 42.107 17 24.798 17
BM25 Robertson and Zaragoza (2009) - 26.211 18 34.155 18 22.582 18
Spearman Rank Correlation Coefficient (P-value) - 0.8204 (3e-5) 0.7028 (1e-3)
@@ 5.1 Consistency Analysis (RQ1)
bge-en-icl 7.11B 71.24 2 61.67 2 53.60 2 72.62 3
bge-en-icl-e5data 7.11B 64.67 10 59.59 6 54.46 1 73.43 2
@@
gte-large-en-v1.5 434M 65.39 9 57.91 9 46.251 11 60.71 14
multilingual-e5-large-instruct 560M 64.41 11 52.47 12 45.39 12 63.96 12
bge-large-en-v1.5 335M 64.23 12 54.29 11 44.91 13 61.86 13
e5-large-v2 335M 62.20 13 50.56 14 46.253 10 66.16 8
multilingual-e5-large 560M 60.89 14 51.40 13 43.91 14 65.71 10
@@
Table 5: Comparison of the performance of 15 IR models on AIR-Bench and MTEB/BEIR. The results on MTEB/BEIR are directly taken from the MTEB leaderboard. For detailed information of the models appearing in the table, please refer to Table 15. The detailed results for each dataset in AIR-Bench are available in Appendix F.
@@
To examine how well G-MSMARCO aligns with R-MSMARCO, we evaluate 18 IR models on R-MSMARCO and G-MSMARCO using nDCG@10, and compute the Spearman rank correlation coefficient Spearman (1961) between their rankings on R-MSMARCO and G-MSMARCO as the consistency metric.
Main Results. As shown in Table 4, the Spearman rank correlation coefficient is 0.8204 with a p-value of 3e-5, indicating that the LLM-generated testing data aligns well with the human-labeled testing data. Overall, each model achieves higher nDCG@10 on G-MSMARCO than on R-MSMARCO. This can be largely attributed to more comprehensive quality control strategy of AIR-Bench, which results in more positives for each query (see Table 6).
Ablation of Quality Control. To demonstrate the necessity of the quality control stage in the data generation pipeline of AIR-Bench, we also evaluate the consistency between R-MSMARCO and G-MSMARCO generated without quality control. As shown in Table 4, the correlation coefficient shows a significant degradation (0.8204 $\rightarrow$ 0.7028). Besides, the nDCG@10 of each model on G-MSMARCO without quality control also has a huge drop, due to some low-quality queries and very limited positives (see Table 6, there are 1,110 low-quality queries and only 7,429 positives). Therefore, quality control stage is necessary to ensure the data generation pipeline a reliable data generation pipeline.
@@
Figure 3: Robustness analysis of the consistency between the LLM-generated testing data and the human-labeled testing data. The mean correlation coefficient is 0.8067 with a mean p-value of 8e-5 across 30 simulated generation processes.
@@ 5.3 Distinguishing Models (RQ3)
Setup. We fine-tune mContriever [^5] https://huggingface.co/facebook/mcontriever-msmarco Izacard et al. (2021), using domain-specific training datasets and compare the model’s performance on the corresponding datasets in AIR-Bench before and after fine-tuning. Specifically, we fine-tune [^6] The learning rate is $2\times 10^{-4}$, the warmup ratio is 0.1, and the weight decay is 0.01. The training process takes around a hundred steps with a total batch size of 64 on 8 A800 GPUs. mContriever with FlagEmbedding tool [^7] https://github.com/FlagOpen/FlagEmbedding to enhance its domain-specific capabilities. The domain-specific training data used for fine-tuning is independent of the corresponding testing data in AIR-Bench.
@@ F.1 Distinguishing Models
Domain. We evaluate three kinds of embedding models with the same model size (large-size), and compare their performances in each domain on AIR-Bench. As shown in Figure 5(c) and Figure 5(d), regardless of whether the task is QA or Long-Doc and whether the datasets are only in English or multilingual, no model is able to achieve the best performance on all domains. For more results, please refer to Appendix F.
Language. We evaluate three kinds of embedding models with the same model size (large-size) on the multilingual datasets of AIR-Bench, and compare their performance on the datasets of each language. As shown in Figure 5(e), we also observe that no model is able to achieve the best performance on all languages. For more results, please refer to Appendix F.
@@ F.2 Detailed Evaluation Results
Figure 5: AIR-Bench can distinguish models in different dimensions, including model dimension, domain dimension, and language dimension. For detailed information of the models appearing in this figure, please refer to Table 15. The detailed metric value and additional results on other model size are all available in Appendix F.
@@
jina-embeddings-v2-base-en Günther et al. (2023) 137M https://huggingface.co/jinaai/jina-embeddings-v2-base-en

### v3

@@ 4 Diversity Analysis
multilingual-e5-small Wang et al. (2024) 118M 42.253 11 47.989 14 28.579 15
simlm-base-msmarco-finetuned Wang et al. (2022a) 110M 41.675 12 48.102 13 30.548 9
jina-embeddings-v3 Sturua et al. (2024) 572M 39.787 13 51.098 12 30.297 12
bge-m3 Chen et al. (2024b) 568M 39.565 14 54.404 6 33.286 3
contriever-msmarco Izacard et al. (2022) 109M 36.570 15 47.127 15 29.231 14
msmarco-roberta-base-ance-firstp Xiong et al. (2021) 125M 33.637 16 42.107 16 24.798 16
BM25 Robertson and Zaragoza (2009) - 26.211 17 34.155 17 22.582 17
Spearman Rank Correlation Coefficient (P-value) - 0.8211 (5e-5) 0.6912 (2e-3)
@@ 5.1 Consistency Analysis (RQ1)
bge-en-icl (zero-shot) 7.11B 71.24 2 61.67 2 53.60 2 72.62 3
bge-en-icl-e5data (zero-shot) 7.11B 64.67 11 59.59 6 54.46 1 73.43 2
@@
jina-embeddings-v3 572M 65.51 9 53.88 12 45.07 13 61.50 13
gte-large-en-v1.5 434M 65.39 10 57.91 9 46.251 11 60.71 14
multilingual-e5-large-instruct 560M 64.41 12 52.47 13 45.39 12 63.96 11
bge-large-en-v1.5 335M 64.23 13 54.29 11 44.91 14 61.86 12
e5-large-v2 335M 62.20 14 50.56 14 46.253 10 66.16 8
@@
Table 5: Comparison of the performance of 15 IR models on AIR-Bench and MTEB/BEIR. The results on MTEB/BEIR are directly taken from the MTEB leaderboard. For detailed information of the models appearing in the table, please refer to Table 15. The detailed results for each dataset in AIR-Bench are available in Appendix F.2.
@@
To examine how well G-MSMARCO aligns with R-MSMARCO, we evaluate 17 IR models on R-MSMARCO and G-MSMARCO using nDCG@10, and compute the Spearman rank correlation coefficient Spearman (1961) between their rankings on R-MSMARCO and G-MSMARCO as the consistency metric.
Main Results. As shown in Table 4, the Spearman rank correlation coefficient is 0.8211 with a p-value of 5e-5, indicating that the LLM-generated testing data aligns well with the human-labeled testing data. Overall, each model achieves higher nDCG@10 on G-MSMARCO than on R-MSMARCO. This can be largely attributed to more comprehensive quality control strategy of AIR-Bench, which results in more positives for each query (see Table 6).
Ablation of Quality Control. To demonstrate the necessity of the quality control stage in the data generation pipeline of AIR-Bench, we also evaluate the consistency between R-MSMARCO and G-MSMARCO generated without quality control. As shown in Table 4, the correlation coefficient shows a significant degradation (0.8211 $\rightarrow$ 0.6912). Besides, the nDCG@10 of each model on G-MSMARCO without quality control also has a huge drop, due to some low-quality queries and very limited positives (see Table 6, there are 1,110 low-quality queries and only 7,429 positives). Therefore, quality control stage is necessary to ensure the data generation pipeline a reliable data generation pipeline.
@@
Figure 3: Robustness analysis of the consistency between the LLM-generated testing data and the human-labeled testing data. The mean correlation coefficient is 0.8031 with a mean p-value of 1e-4 across 30 simulated generation processes.
@@ 5.3 Distinguishing Models (RQ3)
Setup. We fine-tune mContriever [^5] https://huggingface.co/facebook/mcontriever-msmarco Izacard et al. (2021) using domain-specific training datasets, and compare the model’s performance on the corresponding datasets in AIR-Bench before and after fine-tuning. Specifically, we fine-tune [^6] The learning rate is $2\times 10^{-4}$, the warmup ratio is 0.1, and the weight decay is 0.01. The training process takes around a hundred steps with a total batch size of 64 on 8 A800 GPUs. mContriever with FlagEmbedding tool [^7] https://github.com/FlagOpen/FlagEmbedding to enhance its domain-specific capabilities. The domain-specific training data used for fine-tuning is independent of the corresponding testing data in AIR-Bench.
@@ F.1 Distinguishing Models
Domain. We evaluate three kinds of embedding models with the same model size (large-size), and compare their performances in each domain on AIR-Bench. As shown in Figure 5(c) and Figure 5(d), regardless of whether the task is QA or Long-Doc and whether the datasets are only in English or multilingual, no model is able to achieve the best performance on all domains.
Language. We evaluate three kinds of embedding models with the same model size (large-size) on the multilingual datasets of AIR-Bench, and compare their performance on the datasets of each language. As shown in Figure 5(e), we also observe that no model is able to achieve the best performance on all languages.
@@ F.2 Detailed Evaluation Results
Figure 5: AIR-Bench can distinguish models in different dimensions, including model dimension, domain dimension, and language dimension. For detailed information of the models appearing in this figure, please refer to Table 15. The detailed metric value and additional results on other model size are all available in Appendix F.2.